# Enhancing Large Language Models for Text-to-Testcase Generation

### v1

@@ 1. RQ1
    How effective is the proposed approach for the Text-to-Test generation task?
@@ 2. RQ2
    What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
@@ 3. RQ3
    What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
  4. RQ4
@@ 4.2 Evaluation Dataset
To mitigate the risk of data leakage between model fine-tuning and model evaluation, we decided to curate a new benchmark dataset for evaluating the LLMs for the text-to-testcase generation task. Table 1 provides a summary of the studied five open-source Java projects namely, Apache Commons Lang (Lang, 2022), JFreeChart (JFreeChart, 2022), Apache Common CLI (CLI, 2022), Apache Common CSV (CSV, 2022), Google Gson (Gson, 2019). We selected these five projects due to the following reasons. We chose open-source software projects to support the reproducibility and replicability of our work, allowing researchers and practitioners to reproduce experiments and validate our results. We chose diverse domains and varying sizes of projects to ensure that the results are not bound to specific contexts.
@@ 4.3 Baseline Comparisons
To fairly evaluate our proposed approach, we chose the other three competitive LLMs (i.e., GPT-3.5, Bloom and CodeT5) as a baseline comparison. We choose the following three LLMs based on the various model size, types of model architectures, open/closed source, and fit in our GPU infrastructure. Table 2 shows the analysis of the characteristics of the baseline models.
Model Model Size Architecture Open/Closed Source Fit the GPU?
GPT-3.5 175B Decoder only Closed Source No
Bloom 560M Decoder only Open Source Yes
CodeT5 770M Encoder-Decoder Open Source Yes
Table 2: Comparison of Different Models
To investigate the benefits of the fine-tuning and prompting steps for our approach, we conduct an ablation study by selectively removing or ablating one of the components and observing the impact on the overall performance of our approach. Thus, we conducted a comprehensive experiment of the four variations (i.e., With and Without Finetune, and Basic and Advanced Prompt) on three LLMs (i.e., GPT-3.5-turbo, Bloom, CodeT5), totaling 12 models. Since we observed that the fine-tuning and the prompting for Bloom and Code T5 perform best, for the remainder of the paper, we will focus on the comparison of our approach (GPT-3.5-turbo FineTune + Imprv.P), i.e., the prompting-based fine-tuning GPT-3.5-turbo model with the following baseline.
  • 1.
    GPT-3.5-turbo (FineTune + Basic.P)
  • 2.
    GPT-3.5-turbo (NoFineTune + Imprv.P)
  • 3.
    GPT-3.5-turbo (NoFineTune + Basic.P)
  • 4.
    Bloom (Fine-Tune + Imprv.P)
  • 5.
    CodeT5 (Fine-Tune + Imprv.P)
@@
We fine-tune the two models, namely CodeT5 (Wang et al., 2021) and Bloom (Luccioni et al., 2022), using the following settings. We run our experiment on two NVIDIA GeForce RTX 3090 GPUs with 24 GB vRAM, an Intel(R) Core(TM) i9-9980XE CPU@ 3.00GHz with 36 core processors, and 64G RAM. The hyperparameter settings for fine-tuning are reported in Table 3.
@@
Table 3: Hyper-parameters and their values for CodeT5 and Bloom
Figure 4: (RQ1)An example of test cases generated by GPT-3.5-turbo, and other baseline approaches
@@ 4.5 Evaluation Measures
The LLMs take ”Improved prompt and text” as input and the focal method as ground truth for evaluating the generated test cases (Figure 3). We evaluate the quality of the test cases generated by LLMs based on the following three aspects:
  • 1.
@@
    Requirements Alignment measures the percentage of generated test cases that match the specified requirement. Since the ground truth code represents the correct implementation of the method requirement, any test cases that execute successfully against the ground truth code are considered correct (i.e., they correctly align with the given requirement and invoke the appropriate functions to be tested with the correct test inputs and expected outputs). Therefore, we define that a test case is aligned with a given requirement if it can be executed with the ground truth code. To assess this, we execute the combination of the ground-truth code and the generated test cases.
  • 3.
@@
    Code Coverage measures the completeness of the generated test cases for a given description. Ideally, in addition to aligning with the requirements, the test cases must be completed (i.e. executing all lines of the ground-truth code solution). Thus, we measure the line-level code coverage of the generated test cases using the Jacoco (Library, 2019).

### v2

@@ 1. RQ1
    How effective is the proposed approach for the Text-to-Testcase generation task?
@@ 2. RQ2
    What is the effect of the fine-tuning and prompting step for the LLM-based text-to-testcase generation task?
@@ 3.2 Evaluation Dataset
To mitigate the risk of data leakage between model fine-tuning and model evaluation, we decided to curate a new benchmark dataset for evaluating the LLMs for the text-to-testcase generation task. Table 1 provides a summary of the studied five open-source Java projects namely, Apache Commons Lang [25], JFreeChart [26], Apache Common CLI [27], Apache Common CSV [28], Google Gson [29]. We selected these five projects due to the following reasons. We chose open-source software projects to support the reproducibility and replicability of our work, allowing researchers and practitioners to reproduce experiments and validate our results. We chose diverse domains and varying sizes of projects to ensure that the results are not bound to specific contexts.
@@ 3.3 Baseline Comparisons
To fairly evaluate our proposed approach, we chose the other eight LLMs (i.e., GPT-3.5, Incoder, Starcoder, CodeT5, Bloom, CodeGemma, CodeLlama, and Gemini) as a baseline comparison. We choose the following eight LLMs based on the various model size, types of model architectures, open/closed source, and fit in our GPU infrastructure. Table 2 shows the analysis of the characteristics of the baseline models.
Table 2: Baseline Models.
Model Model Size Open/Closed Source
GPT-3.5 NA Closed Source
Bloom 560M Open Source
CodeT5 770M Open Source
Gemini NA Closed Source
Starcoder 137M Open Source
Incoder 6.7M Open Source
CodeLlama 7B Open Source
CodeGemma 7B Open Source
To investigate the benefits of the fine-tuning and prompting steps for our approach, we conduct an ablation study by selectively removing or ablating one of the components and observing the impact on the overall performance of our approach. Thus, we conducted a comprehensive experiment of the four variations (i.e., With and Without Finetune, and Basic and Advanced Prompt) on eight LLMs (i.e., GPT-3.5-turbo, Bloom, CodeT5, Gemini, Starcoder, Incoder, CodeLlama, CodeGemma), totaling 32 models.
@@
We fine-tune the two models, using the following settings. We run our experiment on two NVIDIA GeForce RTX 3090 GPUs with 24 GB vRAM, an Intel(R) Core(TM) i9-9980XE CPU@ 3.00GHz with 36 core processors, and 64G RAM. The hyperparameter settings for fine-tuning are reported in Table 3.
Table 3: Hyper-parameters and their values.
@@ 3.5 Evaluation Measures
The LLMs take ”Improved prompt and text” as input and the focal method as ground truth for evaluating the generated test cases (Figure 3). We evaluate the quality of the test cases generated by LLMs based on the following four aspects:
@@
  • Requirements Alignment measures the percentage of generated test cases that match the specified requirement. Only the syntactically correct test case are considered for the requirement alignment. Since the ground truth code represents the correct implementation of the method requirement, any test cases that execute successfully against the ground truth code are considered correct (i.e., they correctly align with the given requirement and invoke the appropriate functions to be tested with the correct test inputs and expected outputs). Therefore, we define that a test case is aligned with a given requirement if it can be executed with the ground truth code. To assess this, we execute the combination of the ground-truth code and the generated test cases.
@@
  • Code Coverage measures the completeness of the generated test cases for a given description. Ideally, in addition to aligning with the requirements, the test cases must be completed (i.e. executing all lines of the ground-truth code solution). Thus, we measure the line-level code coverage of the generated test cases using the Jacoco [55].
  • Mutation Score measures the effectiveness of the generated test cases. High-quality test cases should be capable of detecting bugs when the code undergoes slight modifications. Mutation testing introduces intentional faults (mutations) into the code to determine if the existing test cases can identify these faults. The mutation score is then calculated as the percentage of mutations detected (killed) out of the total number of mutations introduced. The mutation operators are used to systematically create variants of the program under test, each differing by only a small syntactic change. These operators include Arithmetic Operator Replacement (AOR), Logical Operator Replacement (LOR), Shift Operator Replacement (SOR), Relational Operator Replacement (COR), Increment Operator Replacement (ROR), Unary Operator Replacement (ORU), Literal Value Replacement (LVR), and Statement Deletion (STD). The test cases are executed on each mutant; if a test cases that passes on the original program fails on the mutant, then the mutant is killed. To analyse the mutation score, we use the PIT Test [56] tool to measure the mutation score for our JUnit test cases. We applied all mutation operators: AOR, LOR, SOR, COR, ROR, ORU, LVR, STD. A high mutation score indicates that the generated test case is effective in detecting faults introduced by the mutations.
4 Enhancing Large Language Models for Text-To-Testcase Generation
Problem Formulation. In this paper, we propose a text-to-testcase generation approach based on a fine-tuned LLM with an effective prompting design. In particular, we formulate the text-to-testcase as a generation task using a fine-tuned LLM (GPT-3.5-turbo). The model has been fine-tuned using a large dataset that was specifically mined and curated from various open-source Java projects. During the model fine-tuning, our LLM aims to learn the mapping between the method description (text), which is the comments or documentation that describe the purpose, functionality, and usage of a method, and the corresponding test case (testcase) which is a test method that includes a set of assert statements designed to verify whether the implemented method (method) is aligned with the given requirement or not (i.e., functionally correct). Once fine-tuned, during the model inference phase, an unseen method description is used as input into the fine-tuned LLM to generate the corresponding test cases without requiring the implemented method (method), which differs from most existing test case generation approaches. Our approach consists of three phases: (1) model fine-tuning, (2) prompt design, and (3) model inference. Below, we describe the details of our approach (Figure 1).
4.1 Model Fine-tuning
While LLMs demonstrate impressive capabilities in understanding and generating natural language [37], they may lack the nuanced contextual understanding required for specialized applications, e.g., the text-to-testcase generation task. The fine-tuning of LLMs is essential to enhance their performance and adaptability to our text-to-testcase generation task. Thus, it is important to have a representative and well-labeled dataset for the specific domain to achieve effective fine-tuning results. However, the existing test case generation datasets are not suitable for our task. For example, Tufano et al. [57] introduced the Method2Test dataset, which is a pair of <<<Method,Testcase>>>, but lacking the corresponding method description that is required for our approach. Thus, we need to carefully curate the fine-tuning dataset.
4.1.1 Dataset Curation
For the fine-tuning of the model, we started from the Methods2Test dataset [57]. This dataset consists of 780K pairs of Java focal methods and their corresponding test cases, collected from 91K open-source Java projects. Since the method description is unavailable, we extended the Methods2Test dataset. Starting from the 91K open-source Java projects, we cloned each project from the provided GitHub URL and performed the following steps:
(Step 1) Collecting Code Structural Information. The structural information of the code, such as method and class names are very important as we used to identify the relationship between the focal methods (i.e., methods to be tested) and their corresponding test methods. To collect such information, we used the tree-sitter parser [58] to analyze the structure of source code within the given project based on the detailed abstract syntax tree (AST) representation of source code. With this step, we are able to identify the list of method and class names within the project.
(Step 4) Identify Test Methods, Focal Methods. We identify the test methods for each test class based on the @Test annotation. Similar to Step 3, we apply the method name matching heuristic (i.e., removing Test prefix or suffix in the method name) to identify the corresponding focal method and its method description.
(Step 4) Identify Method Descriptions. We utilized the Java parser to analyze the source code and identify the definitions of methods.
Documentation Extraction: We extracted the accompanying documentation comments, including Javadoc and inline comments, to gather details about each method. This process involved capturing the Javadoc comments placed above methods, which provide a summary of their intended use and behavior. We also extracted the inline comments embedded within the method bodies that offer additional context and explanations. We then aggregated all relevant comments for each method to create a comprehensive description.
With these steps, we construct the dataset <<<Text,Testcase,Method>>> in the format of triplets. The dataset consists of 163k pairs of JUnit test cases and corresponding method description. Our dataset is derived from real-world systems and includes a variety of data characteristics that are typical of development environments, such as incomplete method descriptions and noisy input. Specifically, the dataset consists of method descriptions, comprising 109k pairs with 67% of the total. Inline comments make up 13k pairs, representing 8%, and combinations of method descriptions and inline comments account for 40k pairs with 25% of the dataset. This diversity ensures that our model is evaluated against different textual formats and challenges, including handling missing information. This comprehensive evaluation is crucial for ensuring the practical viability and robustness of your approach in real-world applications. We split the dataset into two sets, i.e., training set (60% - 97,800 pairs), validation set (20% - 32,600 pairs) and test set (20% - 32,600 pairs). We used our validation data set with 32k pairs of <<<Text,Testcase>>> for our finetuning process We note that the methods will not be used for model fine-tuning like prior work, but are only used for evaluating test cases.
Our dataset, used for fine-tuning the language model, comprises software methods with a broad spectrum of characteristics: method inline descriptions range from 16 to 1083 characters (average 39.54), method description vary between 39 and 1672 characters (average 303.52). This collection includes only the method type “public” drawn from various open-source projects to ensure a wide representation of programming styles and complexities. Additionally, the data set is designed to mitigate potential biases by including methods from multiple domains and programming paradigms, with the aim of providing a comprehensive foundation for model training.
4.1.2 Fine-Tuning
When performing model fine-tuning, each training example generally consists of a single input (i.e., method description) example and its associated output (i.e., test case) without giving detailed instructions or including multiple examples in the same prompt normally used for the base models (GPT-3.5). Following the OpenAI’s legacy fine-tuning best practices [^1] https://platform.openai.com/docs/guides/legacy-fine-tuning, each training example typically consists of a single input example and its associated output, which is typically constructed in the following format: {"input": "<Text>", "output": "<test case>"}. We extract method descriptions and corresponding test cases into a list of prompts and completions for fine-tuning. We initiate a fine-tuning job using OpenAI’s CreateFineTuneJob.create method, specifying the training data and model to fine-tune (GPT-3.5-turbo). We use ChatCompletion.create to generate test cases with the fine-tuned model, providing: The model engine ID from the fine-tuning job, the input is the method description prompt. The output is completion of the generated test case.The template includes the {"prompt": "<Text>", "completion": "<test case>". The input is a single input-output pair. We used the curated dataset constructed in Section 4.1.1 for model fine-tuning (the pairs of <text, test case>). The cost for fine-tuning is \$0.0080 per 1,000 tokens, totaling \$68.7 for fine-tuning the GPT-3.5 model.
Once completed, we use the fine-tuned model for generating test cases in the next step.
Figure 2: Structure of our Improved prompt and Basic prompt.
4.2 Model Inference
4.2.1 Test Case Generation
For test case generation, the most important step was to develop a prompt that works well, as the prompts provide context and guidance to the model. Without a (clear) prompt, the model would not know what information or response is expected. The process of formulating prompts for software engineering tasks is inherently subjective, and there are no established empirical guidelines for optimizing prompt effectiveness and robustness. We conducted several experiments involving various prompt combinations, ranging from basic to customized prompts, to address concerns associated with prompt optimization and devised our improved prompt. We present two prompts below. A basic prompt (our starting point) and an improved prompt (after several iterations). We explain each prompting technique below.
Basic prompt For basic prompt, we first design the prompt template as presented in Figure 2 following the OpenAI guidelines  [^2] https://help.openai.com/en/articles/6654000-best-practices-for-promptengineering-with-openai-api  [^3] https://platform.openai.com/docs/guides/prompt-engineering/strategywrite-clear-instructions to ensure that the structure of the prompt is suitable for GPT-3.5. The prompt template consists of the following components: an instruction and an input (i.e., basic instruction and a text description). Finally, we use the LLMs to generate test case from the created prompts. The structure of the basic prompt is as follows:
"/* Write a junit test case /**Description **/"
Improved prompt For Improved prompt, we first design the prompt template as presented in Figure 2. The prompting technique employed is One-shot prompting [42] which involves instructing the model with a single, detailed example or instruction to generate a specific output, such as a JUnit test method. Similar to basic prompt, we follow the guidelines of OpenAI when designing the prompt template. The prompt template consists of the following components: demonstration examples, instruction, and input (that is, the test case template and target class and method name). It is the iterated version based on multistep optimizations. We tested the prompt in the first step on a few samples from a small validation set (not from our curated or evaluation set to avoid data leakage), manually analyzed the generation performance and optimized the generation results by providing ChatGPT with a series of new prompts.
The approach to refining the prompt design for test case generation based on Test-Driven Development (TDD) principles [1] can be aligned with the three core TDD steps: Write a test, run it, and do it right. These principles were applied to disign the improved prompt. We designed the prompt to instruct the LLM with a more specific context. It emphasizes the importance of proper and relevant assertion statements, avoiding repetition, following a structured test case format, and adding context details for a more structured test case. A detailed breakdown of the prompt template components and their intended purpose is as follows:
  • Introductory Comment:(Write a Test) In TDD, the first step is to write a test that fails initially because the functionality it tests does not yet exist. This prompt design, sets the expectations for the model, defining the scope and focus of the test case it needs to generate. Instructs the model to focus on creating a test case, highlighting the need for relevant assert statements and avoiding repetition.
  • Class and Method Name:(Make it Run) After writing a failing test, the next step in TDD is to write enough code stubs to pass the test case. For the prompt design we use the ”Class and Method Name” as our stub, directly guides the LLM to generate a test case relevant to the specific target class and method name for which the test is to be written, ensuring the generated test is relevant.
  • Method Description: The ”Method Description” in the prompt plays a crucial role here, offering the LLM detailed context about what the method does, which enables it to generate more accurate and contextually relevant test assertions. Provides a description of the method, giving context to generate a test case.
4.2.2 Post-Processing
When extracting tests from LLM’s response, syntactic errors may occur because the tokens limit is exceeded, resulting in truncated tests. To address this challenge, we introduce an automated verification approach in order to verify and correct the naming consistency and invalid test signatures. We thus perform a three-part post-processing on the generated test cases.
  1. a)
  2. b)
  3. c)
    Correcting Incomplete Parentheses: We noticed that some of the generated test cases contain incomplete parentheses, such as (or) is incomplete. To address this issue, we perform corrections on these long sequence test cases, ensuring that all parentheses are correctly balanced.
Table 4: (RQ1)The experimental results of the ablation study with eight different models ( Incoder, Starcoder, CodeT5, Bloom, CodeGemma, CodeLlama, Gemini and GPT-3.5).
Evaluation Measures
Project Model Syntax Correctness Requirement Alignment Code Coverage Mutation Score
Chart FT+I.P FT+B.P NoFT+I.P NoFT+B.P FT+I.P FT+B.P NoFT+I.P NoFT+B.P FT+I.P FT+B.P NoFT+I.P NoFT+B.P FT+I.P FT+B.P NoFT+I.P NoFT+B.P
Incoder 1.7 1.14 0.65 0.18 1.48 0.99 0.56 0.15 1.29 0.86 0.49 0.13 0.37 0.25 0.14 0.04
Starcoder 2.4 1.61 0.92 0.25 2.09 1.4 0.8 0.22 1.82 1.22 0.7 0.19 0.53 0.35 0.2 0.06
CodeT5 8.32 7.24 4.13 1.12 3.56 3.1 1.77 0.48 3.03 2.64 1.5 0.41 0.88 0.77 0.44 0.12
Bloom 26.02 22.64 12.9 3.48 22.94 19.96 11.38 3.07 15.93 13.86 7.9 2.13 4.62 4.02 2.29 0.62
CodeGemma 49.7 33.3 18.98 5.12 43.24 28.97 16.51 4.46 37.62 25.21 14.37 3.88 10.91 7.31 4.17 1.13
CodeLlama 56.79 38.05 21.69 5.86 49.41 33.1 18.87 5.09 42.99 28.8 16.42 4.43 12.47 8.35 4.76 1.28
Gemini 74.8 50.12 28.57 7.71 65.08 43.6 24.85 6.71 56.62 37.94 21.63 5.84 16.42 11 6.27 1.69
GPT-3.5 82.4 46.09 26.48 17.4 73.7 35.17 25.19 13.91 69.6 31.43 21.3 12.91 20.18 9.11 6.18 3.74
Cli Incoder 1.67 1.12 0.64 0.17 1.45 0.97 0.55 0.15 1.26 0.84 0.48 0.13 0.37 0.24 0.14 0.04
Starcoder 2.35 1.57 0.89 0.24 2.04 1.37 0.78 0.21 1.77 1.19 0.68 0.18 0.51 0.35 0.2 0.05
CodeT5 7.3 6.35 3.62 0.98 1.9 1.65 0.94 0.25 1.74 1.51 0.86 0.23 0.5 0.44 0.25 0.07
Bloom 16.3 14.18 8.08 2.18 11.89 10.34 5.89 1.59 10.44 9.08 5.18 1.4 3.03 2.63 1.5 0.41
CodeGemma 48.71 32.64 18.6 5.02 42.38 28.39 16.18 4.37 36.87 24.7 14.08 3.8 10.69 7.16 4.08 1.1
CodeLlama 55.65 37.29 21.26 5.74 48.42 32.44 18.49 4.99 42.13 28.23 16.09 4.34 12.22 8.19 4.67 1.26
Gemini 67 51.59 29.41 7.94 66.99 44.88 25.58 6.91 58.28 39.05 22.26 6.01 16.9 11.32 6.46 1.74
GPT-3.5 73.7 37.8 26.08 11.08 62.5 26.05 15.05 9.81 59.5 23.04 12.78 8.09 17.26 6.68 3.71 2.35
Csv Incoder 1.09 0.73 0.42 0.11 0.95 0.64 0.36 0.1 0.83 0.56 0.32 0.09 0.24 0.16 0.09 0.03
Starcoder 2.67 1.79 1.02 0.28 2.32 1.55 0.88 0.24 2.02 1.35 0.77 0.21 0.59 0.39 0.22 0.06
CodeT5 7.05 6.13 3.49 0.94 2.07 1.8 1.03 0.28 1.78 1.55 0.88 0.24 0.52 0.45 0.26 0.07
Bloom 17.57 15.29 8.72 2.35 12.07 10.5 5.99 1.62 11.98 10.42 5.94 1.6 3.47 3.02 1.72 0.46
CodeGemma 44.78 30 17.1 4.62 38.96 26.1 14.88 4.02 33.9 22.71 12.94 3.49 9.83 6.59 3.75 1.01
CodeLlama 50.5 33.84 19.29 5.21 43.94 29.44 16.78 4.53 38.23 25.61 14.6 3.94 11.09 7.43 4.23 1.14
Gemini 71 47.57 27.11 7.32 61.77 41.39 23.59 6.37 53.74 36.01 20.53 5.54 15.58 10.44 5.95 1.61
GPT-3.5 77.8 38.56 28.78 10.7 64.1 27.06 18.01 10.14 61.1 24.16 14.16 9.92 17.72 7.01 4.11 2.88
Gson Incoder 1.99 1.33 0.76 0.21 1.73 1.16 0.66 0.18 1.51 1.01 0.58 0.16 0.44 0.29 0.17 0.05
Starcoder 3.23 2.16 1.23 0.33 2.81 1.88 1.07 0.29 2.44 1.63 0.93 0.25 0.71 0.47 0.27 0.07
CodeT5 7.31 6.36 3.63 0.98 3.14 2.73 1.56 0.42 2.67 2.32 1.32 0.36 0.77 0.67 0.38 0.1
Bloom 19.73 17.17 9.79 2.64 15.14 13.17 7.51 2.03 13.67 11.89 6.78 1.83 3.96 3.45 1.97 0.53
CodeGemma 49.08 32.88 18.74 5.06 42.7 28.61 16.31 4.4 37.15 24.89 14.19 3.83 10.77 7.22 4.12 1.11
CodeLlama 55.89 37.45 21.35 5.76 48.62 32.58 18.57 5.01 42.3 28.34 16.15 4.36 12.27 8.22 4.68 1.26
Gemini 70.9 47.5 27.08 7.31 61.68 41.33 23.56 6.36 53.66 35.95 20.49 5.53 15.56 10.43 5.94 1.6
GPT-3.5 80.5 39.89 29.84 9.91 70.9 28.1 20.01 9.82 70.8 24.49 14.89 8.72 20.53 7.1 4.32 2.53
Lang Incoder 1.06 0.71 0.4 0.11 0.92 0.62 0.35 0.09 0.8 0.54 0.31 0.08 0.23 0.16 0.09 0.02
Starcoder 2.98 2 1.14 0.31 2.59 1.74 0.99 0.27 2.25 1.51 0.86 0.23 0.65 0.44 0.25 0.07
CodeT5 6.9 6 3.42 0.92 1.56 1.36 0.78 0.21 1.05 0.91 0.52 0.14 0.3 0.26 0.15 0.04
Bloom 23.99 20.87 11.9 3.21 19.86 17.28 9.85 2.66 13.05 11.35 6.47 1.75 3.78 3.29 1.88 0.51
CodeGemma 41.45 27.77 15.83 4.27 36.06 24.16 13.77 3.72 31.37 21.02 11.98 3.23 9.1 6.1 3.47 0.94
CodeLlama 49.01 32.84 18.72 5.05 42.64 28.57 16.28 4.4 37.1 24.86 14.17 3.83 10.76 7.21 4.11 1.11
Gemini 68.05 45.59 25.99 7.02 59.2 39.66 22.61 6.1 51.5 34.51 19.67 5.31 14.94 10.01 5.7 1.54
GPT-3.5 76.9 45.8 33.8 15.74 63.9 32.4 22.47 11.87 62.9 27.4 17.98 10.95 18.24 7.95 5.21 3.18